40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Tue Mar 15 00:30:46 2011, Zach, Laser, stuff happens, recent gyro plots are LIES 
    Reply  Tue Mar 15 12:34:33 2011, Alastair, Laser, stuff happens, recent gyro plots are LIES 
Message ID: 1353     Entry time: Tue Mar 15 00:30:46 2011     Reply to this: 1354
Author: Zach 
Type: Laser 
Category: stuff happens 
Subject: recent gyro plots are LIES 

 Due to an error somewhere in the front end code, the channel that should have been "PLL_ACT" was actually receiving the "CCW_REFL_DC" signal. All data and plots for noise in the PLL actuation readout are absolutely wrong.

Having made sure that I was monitoring the right damn channel, I observe that the PLL actuation noise is once again the same as the AOM actuation signal. This puts us at ~2 x 10-5 (rad/s)/rHz at our most lenient requirement frequency of 200 mHz, or roughly 2700x where we need to be. I currently have and all enclosures shut and all boost engaged (2 for the primary loop and 1 for the secondary). Changing the number of boosts does not affect the low frequency noise relationship between AOM_ACT and PLL_ACT. The secondary loop has >60 dB of OL gain below 50 Hz, so if changing the boosts does not affect the noise then it means that either 1) the same noise that couples into the AOM signal is coupling in independently into the PLL loop and causing the same noise level despite the suppression from the secondary loop, or b) we don't understand how it really works. I cannot see how differential IO noise is not suppressed by the secondary loop gain as measured in the PLL actuation; this just doesn't make sense.

I'm not really sure what else to do. We have isolated all the optics from the open air, and ensured that the locking loops have a pretty solid amount of gain (the secondary loop gain is not astronomical, but it's enough that we should see something) and yet we can't reduce the noise at the transmitted port. The addition of the vacuum system and custom PDs has amounted to a factor of ~5 better noise. Big whoop.

Also, the chamber pressure is now up to about 1.7 torr (no factors of 10 here).

ELOG V3.1.3-